# Conformal Reconfigurable Intelligent Surfaces: A Cylindrical Geometry Perspective

### v1

@@ I Introduction
Reconfigurable intelligent surfaces (RISs) and metasurfaces have emerged as promising technologies for future wireless communication systems, offering unprecedented control over electromagnetic (EM) wave propagation across the microwave, millimeter-wave, and terahertz domains [Bas2019, Cheng:2022ri, Malevich:2025vl]. By dynamically adjusting the reflection properties of subwavelength elements, these engineered surfaces can tailor the wireless channel in real time, enabling functions such as coverage enhancement, interference suppression, and improved link reliability in environments where infrastructure is sparse, obstructed, or highly dynamic [Liu:201ri].
Owing to their low-power, lightweight, and programmable nature, RISs are particularly suited to dual-use scenarios, encompassing both civilian and tactical applications. Platforms such as unmanned aerial vehicles (UAVs) [Li2021] and connected vehicles [Naaz:2024et] can exploit RIS-assisted communication to improve spectral efficiency and maintain robust connectivity under non-line-of-sight conditions. In dense urban environments, RISs can enhance vehicle-to-everything, UAV-to-infrastructure, and UAV-to-vehicle links, where multipath and shadowing effects often degrade performance [Kisselef:2020ri]. More broadly, the smart radio environment paradigm [DiR2020] envisions RISs as passive yet reconfigurable elements that reshape the wireless channel itself, enabling advanced functions such as multi-hop relaying and wavefront control without relying on traditional active relays. These advantages make RISs attractive for flexible deployment across both mobile platforms and fixed infrastructure, including in disaster recovery, temporary coverage extension, and secure battlefield communications.
@@
In the EM research community, conformal metasurfaces have been extensively explored as versatile platforms for manipulating waves on curved geometries, with applications ranging from optics to microwaves. On the modeling side, advances such as the conformal boundary optics framework have extended generalized sheet transition conditions to free-form surfaces, enabling accurate treatment of curvature-induced phase accumulation and modified boundary conditions [Wu:2018mo]. At optical frequencies, flexible all-dielectric metasurfaces have enabled lensing, cloaking, and wavefront shaping on curved carriers, with experiments showing good tolerance to mechanical deformation and variations in incidence angle [Cheng:2016ad]. Recent reviews on flexible and conformal metasurfaces [Zhou:2024fm] have emphasized practical fabrication aspects, including stretchable substrates, transfer printing, and integration on complex three-dimensional (3D) platforms, while also noting challenges related to alignment, robustness under bending, and long-term reliability.
@@
In the microwave and millimeter-wave domains, conformal metasurfaces have been widely investigated for beam steering and radiation control. Early work demonstrated active cylindrical metasurface lenses that use varactor tuning for wide-angle steering [Li:2019wa]. Later studies showed that passive conformal reflective metasurfaces can enhance beam manipulation and gain in 5G millimeter-wave multi-input multi-output configurations [Malik:2024bs]. Transmissive conformal metasurface antennas implemented on arbitrary cylindrical supports further confirmed the feasibility of practical curved apertures [Liu:2023ct]. At terahertz frequencies, beam steering on flexible curved substrates has been shown to maintain the desired deflection performance across a broad range of bending radii [Shiri:2023tb]. Additional progress includes data-driven, real-time microwave reconfigurability based on adaptive reflective surfaces [Wen:2023rd] and reconfigurable spin-decoupled conformal metasurfaces fabricated on 3D-printed curved substrates [Fu:2025rs]. Recent developments also encompass conformal holographic metasurfaces for scattering-pattern modulation [Zhang:2024am], multifunctional time-modulated conformal designs for radar and spectrum-control applications [Sun:2024af, Sun:2024fc], reconfigurable free-form camouflage metasurfaces [Li:2018rf], and conformal reconfigurable holographic metasurfaces for multifunctional radiation and scattering manipulation [Zhang:2025cr]. Collectively, these results demonstrate significant opportunities offered by conformal platforms while also highlighting practical challenges associated with curvature-dependent performance, fabrication tolerances, and packaging on nonplanar carriers.
@@
From a communication-oriented perspective, the influence of RIS geometry has begun to receive systematic attention [Mizmizi:2023cm]. Comparative analyses of planar, linear, and cylindrical RIS topologies [Cui:2023io] indicate that curvature affects illumination uniformity, effective aperture, and near-field interactions. These effects translate into measurable differences in link budget, signal-to-noise ratio, and outage probability. Such observations underscore the importance of modeling approaches that connect EM behavior, hardware constraints, and communication-level performance metrics.
@@
Within this broader context, conformal RIS platforms emerge as a particularly attractive solution. By following the natural curvature of the supporting structure, conformal RISs enable seamless integration on real-world platforms while introducing additional geometrical degrees of control, which can be exploited to enhance beam steering, improve coverage shaping, and increase spatial diversity [An:2025fi, Mursia:2025t3]. In this sense, curvature itself becomes a functional design parameter, enabling more flexible control of radiation patterns and wavefront shaping than is attainable with conventional planar RIS implementations.
@@ II-B Idealized Scenario: Surface Impedance
To explore the fundamental physical limitations of cylindrical metasurfaces, in analogy with the planar case studied in Ref. [DiRenzo:2022cm], we begin by analyzing an idealized, isotropic scenario. Specifically, we consider a cylinder of radius $R$, characterized by a continuous, scalar surface impedance distribution $Z(\varphi)$, as illustrated in Figure 2a. This model imposes the following boundary condition at $r=R$ [Hoppe:1995ib]:
@@
This problem can be addressed analytically using a Bessel-Fourier expansion. The scattered electric field can be expressed as [Harrington:2001th]
@@
where $\mathcal{H}_{m}^{(2)}(\cdot)$ is the $m$th-order Hankel function of the second kind [Abramowitz:1965ho], and $c_{m}$ are expansion coefficients.
@@
Using the Jacobi-Anger expansion [Abramowitz:1965ho], the corresponding coefficients in Equation (3) can be determined as:
@@
where $\mathcal{J}_{m}(\cdot)$ is the $m$th-order Bessel function of the first kind [Abramowitz:1965ho].
@@
where $\eta_{0}$ is the free-space intrinsic impedance, and the overdot indicates differentiation with respect to the argument. As is well known, the infinite series in Equation (7) can be truncated to a finite number of terms $M\sim k_{0}R$, representing the number of available EM degrees of freedom [Bucci:1989ot].
@@
However, as previously observed in the planar case [DiRenzo:2022cm], this approach generally ensures only global passivity, without guaranteeing local passivity. The resulting impedance profile in Equation (7) is, in general, complex-valued, indicating that different regions of the surface may exhibit gain or loss, even if the overall structure may remain passive on average.
Figure 3 presents representative numerical results. In particular, Figure 3a,b show the normalized surface impedance computed from Equation (7) for a steering angle of $\varphi_{o}=15^{\circ}$. As expected, the surface resistance exhibits both positive and negative values, indicating the coexistence of locally lossy and active (gain) regions. Similar behaviors are observed for other steering angles (not shown for brevity). Figure 3c displays the far-field scattering patterns for steering angles ranging from $15^{\circ}$ to $75^{\circ}$ in $15^{\circ}$ increments. The patterns remain consistent in shape and peak amplitude across the range of steering angles, while maintaining effective sidelobe level (SLL) control. Notably, forward scattering is significantly suppressed, a behavior reminiscent of parity-time-symmetric invisibility cloaking mechanisms [Hao:2019ua], where a tailored balance of gain and loss is used to minimize the total scattering signature.
@@
To ensure local passivity of the structure, an alternative approach can be adopted. This method approximates the cylindrical surface locally by its tangent plane and applies the well-established geometrical-optics (GO) framework developed for planar anomalous reflectors [Sun:2012he]. A detailed derivation is provided in the Appendix. This formulation yields a purely imaginary surface impedance given by
@@
This approach inherits the known limitations of GO-based methods in the synthesis of planar anomalous reflectors [Diaz:2017ft]. Moreover, due to the tangent-plane approximation, its validity is further restricted to configurations with electrically large curvature radii, where $k_{0}R\gg 1$.
@@
This trend is consistent with observations in the planar case [Diaz:2017ft]. In that context, it was shown that globally passive designs, based on a balance of gain and loss, can be effectively synthesized using fully passive structures by carefully engineering strong spatial dispersion. While this approach could, in principle, be extended to cylindrical geometries, its implementation would be considerably more complex. Furthermore, spatially dispersive structures are inherently difficult to reconfigure, which limits their practicality in dynamic RIS applications.
@@ II-C Realistic Scenario: Reconfigurable Digital Elements
Both idealized approaches discussed earlier require surface-impedance profiles with rapid spatial variations and large dynamic ranges, which pose significant challenges for implementation on low-complexity, reconfigurable platforms. To address this, we consider a more practical scenario, illustrated in Figure 2b, in which the cylindrical metasurface is composed of identical elements (“meta-atoms”) with reconfigurable reflection properties [Cui:2014cm].
@@
Our model is based on a physical-optics approximation [Osipov:2017me] and is inspired by the theory of conformal (circular) arrays [Persson:2006th]. Within this framework, the far-field response is approximated as
@@
In principle, the model described above can be extended to more general geometries with variable curvature, as long as the fundamental assumptions of the physical-optics approximation remain valid. In particular, multiple reflections, which are not accounted for in the current formulation, should be negligible for the approximation to hold [Osipov:2017me].
@@
The formulation in Equation (14) constitutes a mixed-integer fractional quadratic programming problem, which is generally challenging to solve efficiently due to its non-convex and combinatorial nature [Lee2011]. When the solution space becomes large, i.e., when $|\Gamma|=L^{N}$ grows, an exhaustive search (ES) quickly becomes computationally infeasible. As a result, alternative approaches such as relaxation techniques and evolutionary algorithms are commonly employed to find near-optimal solutions [Ata2020, Li2021, Lin2020, Yue2023, Che2024].
In our study, we evaluated both genetic algorithms (GAs) [Haupt:2007ga] and low-complexity methods, such as the MPDR beamforming technique [Cap1969]. The latter, detailed in the Appendix, yields a closed-form solution. As an additional low-complexity alternative, we also considered a brute-force spatial sampling and phase quantization of the GO solution given in Equation (8), which likewise results in a closed-form expression.
Note that the choice of beamwidth $\Delta\varphi$ is not entirely arbitrary, as it is physically constrained by the electrical size of the structure, i.e., the number of EM degrees of freedom available [Bucci:1989ot]. Intuitively, smaller structures produce less directional scattering, while larger ones can support narrower beams. As a result, assigning an excessively small beamwidth may lead to an ill-conditioned synthesis problem. As a physically meaningful reference, we adopt the beamwidth of a uniform circular array with the same size and number of elements [Persson:2006th].
@@
As a reconfigurable meta-atom, we adopt the design proposed in Ref. [Yang:2016ap], consisting of a square metallic patch connected to the ground plane through a vertical via controlled by a positive-intrinsic-negative (PIN) diode. The geometry is scaled appropriately for operation at the target frequency. Figure 7 shows the meta-atom structure and its angular response, obtained through full-wave simulations at 3.6 GHz (see the Methods section for mode details). This response is used in the semi-analytical model of Equation (10) to define the local reflection coefficients.
As can be observed in Figure S1, the reflection magnitude remains stable and close to unity across the range of incidence angles, while the phase difference between the two states is exactly $180^{\circ}$ at normal incidence but gradually deteriorates at oblique angles. Consequently, meta-atoms located near the edges of the illuminated region become less effective in switching between states, thereby reducing the number of available degrees of control. Although this particular design is not optimized for angular stability, and improved configurations have been proposed in the literature [Shabanpour:2024eo], the phase difference will inherently vanish as the incidence approaches grazing.
@@ II-E Additional Considerations
The proposed cylindrical RIS is designed for operation at a single frequency where the two meta-atom states exhibit an ideal 180 degree phase difference under normal incidence. To assess the impact of frequency detuning, we carried out additional simulations, with representative results reported in Figure 9. As the operating frequency departs from its design value of 3.6 GHz, the reflection magnitude of the meta-atom in both ON and OFF states under normal incidence (Figure S3a) exhibits only mild variations, whereas the phase contrast between the two states (Figure S3b) degrades more noticeably. As shown in Figure S3c, this mainly affects the sidelobe structure of the scattered field, while the main beam direction and amplitude remain comparatively stable over the considered frequency range. This behavior is consistent with the narrowband nature of the selected meta-atom, which was not designed for dispersion or angular stability. Wider operational bandwidths could be achieved by employing meta-atoms specifically engineered for frequency-stable phase response, such as multi-resonant or dispersion-flattened designs reported in the literature [Simovski:2025bo]. Integrating such elements into the present cylindrical RIS architecture is feasible and represents a natural direction for future developments.
@@
Regarding the influence of element density, the dimensions used in this work are consistent with typical RIS implementations at 3.6 GHz and reflect practical constraints imposed by UAV-scale cylindrical platforms. Although smaller reconfigurable meta-atoms on the order of $\lambda_{0}/5$ have been demonstrated [Dai:2025jc] and could, in principle, increase the number of available degrees of control, this benefit must be balanced against the added hardware complexity and the reduced area available for biasing networks. It is also important to note that beam-pointing accuracy is mainly limited by phase quantization, whereas cell spacing mainly affects sidelobe behavior once near-Nyquist sampling is satisfied. Further reduction in meta-atom size would make diode integration and bias routing more challenging. Within these considerations, the chosen element dimensions provide an effective compromise between practical implementability and beamforming performance for cylindrical RIS deployments.
@@
Finally, in the present study, the polarization sensitivity is essentially dictated by the adopted meta-atom geometry rather than by the cylindrical conformal arrangement itself. In particular, the asymmetric loading associated with the PIN diode breaks the unit-cell symmetry, resulting in a polarization-selective response. The cylindrical geometry mainly affects the local angle of incidence, while the polarization response remains governed by the meta-atom design. It is worth noting that polarization-insensitive metasurfaces have been widely reported (see, e.g., [Li:2024mp]). These approaches could be directly extended to cylindrical RISs to mitigate polarization sensitivity. A systematic investigation of polarization effects in conformal RISs, including cross-polar coupling induced by curvature, will be addressed in future work.

### v2

@@ Abstract
This article has been accepted for publication in a future issue of Wiley Advanced Electronic Materials, but has not been fully edited. Content may change prior to final publication.
Citation information: DOI 10.1002/aelm.202500550, Wiley Advanced Electronic Materials.
@@ I Introduction
Reconfigurable intelligent surfaces (RISs) and metasurfaces have emerged as promising technologies for future wireless communication systems, offering unprecedented control over electromagnetic (EM) wave propagation across the microwave, millimeter-wave, and terahertz domains [5, 10, 32]. By dynamically adjusting the reflection properties of subwavelength elements, these engineered surfaces can tailor the wireless channel in real time, enabling functions such as coverage enhancement, interference suppression, and improved link reliability in environments where infrastructure is sparse, obstructed, or highly dynamic [31].
Owing to their low-power, lightweight, and programmable nature, RISs are particularly suited to dual-use scenarios, encompassing both civilian and tactical applications. Platforms such as unmanned aerial vehicles (UAVs) [28] and connected vehicles [37] can exploit RIS-assisted communication to improve spectral efficiency and maintain robust connectivity under non-line-of-sight conditions. In dense urban environments, RISs can enhance vehicle-to-everything, UAV-to-infrastructure, and UAV-to-vehicle links, where multipath and shadowing effects often degrade performance [22]. More broadly, the smart radio environment paradigm [15] envisions RISs as passive yet reconfigurable elements that reshape the wireless channel itself, enabling advanced functions such as multi-hop relaying and wavefront control without relying on traditional active relays. These advantages make RISs attractive for flexible deployment across both mobile platforms and fixed infrastructure, including in disaster recovery, temporary coverage extension, and secure battlefield communications.
@@
In the EM research community, conformal metasurfaces have been extensively explored as versatile platforms for manipulating waves on curved geometries, with applications ranging from optics to microwaves. On the modeling side, advances such as the conformal boundary optics framework have extended generalized sheet transition conditions to free-form surfaces, enabling accurate treatment of curvature-induced phase accumulation and modified boundary conditions [46]. At optical frequencies, flexible all-dielectric metasurfaces have enabled lensing, cloaking, and wavefront shaping on curved carriers, with experiments showing good tolerance to mechanical deformation and variations in incidence angle [8]. Recent reviews on flexible and conformal metasurfaces [52] have emphasized practical fabrication aspects, including stretchable substrates, transfer printing, and integration on complex three-dimensional (3D) platforms, while also noting challenges related to alignment, robustness under bending, and long-term reliability.
@@
In the microwave and millimeter-wave domains, conformal metasurfaces have been widely investigated for beam steering and radiation control. Early work demonstrated active cylindrical metasurface lenses that use varactor tuning for wide-angle steering [26]. Later studies showed that passive conformal reflective metasurfaces can enhance beam manipulation and gain in 5G millimeter-wave multi-input multi-output configurations [33]. Transmissive conformal metasurface antennas implemented on arbitrary cylindrical supports further confirmed the feasibility of practical curved apertures [30]. At terahertz frequencies, beam steering on flexible curved substrates has been shown to maintain the desired deflection performance across a broad range of bending radii [40]. Additional progress includes data-driven, real-time microwave reconfigurability based on adaptive reflective surfaces [45] and reconfigurable spin-decoupled conformal metasurfaces fabricated on 3D-printed curved substrates [17]. Recent developments also encompass conformal holographic metasurfaces for scattering-pattern modulation [50], multifunctional time-modulated conformal designs for radar and spectrum-control applications [42, 43], reconfigurable free-form camouflage metasurfaces [27], and conformal reconfigurable holographic metasurfaces for multifunctional radiation and scattering manipulation [51]. Collectively, these results demonstrate significant opportunities offered by conformal platforms while also highlighting practical challenges associated with curvature-dependent performance, fabrication tolerances, and packaging on nonplanar carriers.
@@
From a communication-oriented perspective, the influence of RIS geometry has begun to receive systematic attention [35]. Comparative analyses of planar, linear, and cylindrical RIS topologies [12] indicate that curvature affects illumination uniformity, effective aperture, and near-field interactions. These effects translate into measurable differences in link budget, signal-to-noise ratio, and outage probability. Such observations underscore the importance of modeling approaches that connect EM behavior, hardware constraints, and communication-level performance metrics.
@@
Within this broader context, conformal RIS platforms emerge as a particularly attractive solution. By following the natural curvature of the supporting structure, conformal RISs enable seamless integration on real-world platforms while introducing additional geometrical degrees of control, which can be exploited to enhance beam steering, improve coverage shaping, and increase spatial diversity [2, 36]. In this sense, curvature itself becomes a functional design parameter, enabling more flexible control of radiation patterns and wavefront shaping than is attainable with conventional planar RIS implementations.
@@ II-B Idealized Scenario: Surface Impedance
To explore the fundamental physical limitations of cylindrical metasurfaces, in analogy with the planar case studied in Ref. [14], we begin by analyzing an idealized, isotropic scenario. Specifically, we consider a cylinder of radius $R$, characterized by a continuous, scalar surface impedance distribution $Z(\varphi)$, as illustrated in Figure 2a. This model imposes the following boundary condition at $r=R$ [20]:
@@
This problem can be addressed analytically using a Bessel-Fourier expansion. The scattered electric field can be expressed as [18]
@@
where $\mathcal{H}_{m}^{(2)}(\cdot)$ is the $m$th-order Hankel function of the second kind [1], and $c_{m}$ are expansion coefficients.
@@
Using the Jacobi-Anger expansion [1], the corresponding coefficients in Equation (3) can be determined as:
@@
where $\mathcal{J}_{m}(\cdot)$ is the $m$th-order Bessel function of the first kind [1].
@@
where $\eta_{0}$ is the free-space intrinsic impedance, and the overdot indicates differentiation with respect to the argument. As is well known, the infinite series in Equation (7) can be truncated to a finite number of terms $M\sim k_{0}R$, representing the number of available EM degrees of freedom [6].
@@
However, as previously observed in the planar case [14], this approach generally ensures only global passivity, without guaranteeing local passivity. The resulting impedance profile in Equation (7) is, in general, complex-valued, indicating that different regions of the surface may exhibit gain or loss, even if the overall structure may remain passive on average.
Figure 3 presents representative numerical results. In particular, Figure 3a,b show the normalized surface impedance computed from Equation (7) for a steering angle of $\varphi_{o}=15^{\circ}$. As expected, the surface resistance exhibits both positive and negative values, indicating the coexistence of locally lossy and active (gain) regions. Similar behaviors are observed for other steering angles (not shown for brevity). Figure 3c displays the far-field scattering patterns for steering angles ranging from $15^{\circ}$ to $75^{\circ}$ in $15^{\circ}$ increments. The patterns remain consistent in shape and peak amplitude across the range of steering angles, while maintaining effective sidelobe level (SLL) control. Notably, forward scattering is significantly suppressed, a behavior reminiscent of parity-time-symmetric invisibility cloaking mechanisms [25], where a tailored balance of gain and loss is used to minimize the total scattering signature.
@@
To ensure local passivity of the structure, an alternative approach can be adopted. This method approximates the cylindrical surface locally by its tangent plane and applies the well-established geometrical-optics (GO) framework developed for planar anomalous reflectors [44]. A detailed derivation is provided in the Appendix. This formulation yields a purely imaginary surface impedance given by
@@
This approach inherits the known limitations of GO-based methods in the synthesis of planar anomalous reflectors [16]. Moreover, due to the tangent-plane approximation, its validity is further restricted to configurations with electrically large curvature radii, where $k_{0}R\gg 1$.
@@
This trend is consistent with observations in the planar case [16]. In that context, it was shown that globally passive designs, based on a balance of gain and loss, can be effectively synthesized using fully passive structures by carefully engineering strong spatial dispersion. While this approach could, in principle, be extended to cylindrical geometries, its implementation would be considerably more complex. Furthermore, spatially dispersive structures are inherently difficult to reconfigure, which limits their practicality in dynamic RIS applications.
@@ II-C Realistic Scenario: Reconfigurable Digital Elements
Both idealized approaches discussed earlier require surface-impedance profiles with rapid spatial variations and large dynamic ranges, which pose significant challenges for implementation on low-complexity, reconfigurable platforms. To address this, we consider a more practical scenario, illustrated in Figure 2b, in which the cylindrical metasurface is composed of identical elements (“meta-atoms”) with reconfigurable reflection properties [11].
@@
Our model is based on a physical-optics approximation [38] and is inspired by the theory of conformal (circular) arrays [21]. Within this framework, the far-field response is approximated as
@@
In principle, the model described above can be extended to more general geometries with variable curvature, as long as the fundamental assumptions of the physical-optics approximation remain valid. In particular, multiple reflections, which are not accounted for in the current formulation, should be negligible for the approximation to hold [38].
@@
The formulation in Equation (14) constitutes a mixed-integer fractional quadratic programming problem, which is generally challenging to solve efficiently due to its non-convex and combinatorial nature [23]. When the solution space becomes large, i.e., when $|\Gamma|=L^{N}$ grows, an exhaustive search (ES) quickly becomes computationally infeasible. As a result, alternative approaches such as relaxation techniques and evolutionary algorithms are commonly employed to find near-optimal solutions [4, 28, 29, 48, 9].
In our study, we evaluated both genetic algorithms (GAs) [19] and low-complexity methods, such as the MPDR beamforming technique [7]. The latter, detailed in the Appendix, yields a closed-form solution. As an additional low-complexity alternative, we also considered a brute-force spatial sampling and phase quantization of the GO solution given in Equation (8), which likewise results in a closed-form expression.
Note that the choice of beamwidth $\Delta\varphi$ is not entirely arbitrary, as it is physically constrained by the electrical size of the structure, i.e., the number of EM degrees of freedom available [6]. Intuitively, smaller structures produce less directional scattering, while larger ones can support narrower beams. As a result, assigning an excessively small beamwidth may lead to an ill-conditioned synthesis problem. As a physically meaningful reference, we adopt the beamwidth of a uniform circular array with the same size and number of elements [21].
@@
As a reconfigurable meta-atom, we adopt the design proposed in Ref. [47], consisting of a square metallic patch connected to the ground plane through a vertical via controlled by a positive-intrinsic-negative (PIN) diode. The geometry is scaled appropriately for operation at the target frequency. Figure 7 shows the meta-atom structure and its angular response, obtained through full-wave simulations at 3.6 GHz (see the Methods section for mode details). This response is used in the semi-analytical model of Equation (10) to define the local reflection coefficients.
As can be observed in Figure S1, the reflection magnitude remains stable and close to unity across the range of incidence angles, while the phase difference between the two states is exactly $180^{\circ}$ at normal incidence but gradually deteriorates at oblique angles. Consequently, meta-atoms located near the edges of the illuminated region become less effective in switching between states, thereby reducing the number of available degrees of control. Although this particular design is not optimized for angular stability, and improved configurations have been proposed in the literature [39], the phase difference will inherently vanish as the incidence approaches grazing.
@@ II-E Additional Considerations
The proposed cylindrical RIS is designed for operation at a single frequency where the two meta-atom states exhibit an ideal 180 degree phase difference under normal incidence. To assess the impact of frequency detuning, we carried out additional simulations, with representative results reported in Figure 9. As the operating frequency departs from its design value of 3.6 GHz, the reflection magnitude of the meta-atom in both ON and OFF states under normal incidence (Figure S3a) exhibits only mild variations, whereas the phase contrast between the two states (Figure S3b) degrades more noticeably. As shown in Figure S3c, this mainly affects the sidelobe structure of the scattered field, while the main beam direction and amplitude remain comparatively stable over the considered frequency range. This behavior is consistent with the narrowband nature of the selected meta-atom, which was not designed for dispersion or angular stability. Wider operational bandwidths could be achieved by employing meta-atoms specifically engineered for frequency-stable phase response, such as multi-resonant or dispersion-flattened designs reported in the literature [41]. Integrating such elements into the present cylindrical RIS architecture is feasible and represents a natural direction for future developments.
@@
Regarding the influence of element density, the dimensions used in this work are consistent with typical RIS implementations at 3.6 GHz and reflect practical constraints imposed by UAV-scale cylindrical platforms. Although smaller reconfigurable meta-atoms on the order of $\lambda_{0}/5$ have been demonstrated [13] and could, in principle, increase the number of available degrees of control, this benefit must be balanced against the added hardware complexity and the reduced area available for biasing networks. It is also important to note that beam-pointing accuracy is mainly limited by phase quantization, whereas cell spacing mainly affects sidelobe behavior once near-Nyquist sampling is satisfied. Further reduction in meta-atom size would make diode integration and bias routing more challenging. Within these considerations, the chosen element dimensions provide an effective compromise between practical implementability and beamforming performance for cylindrical RIS deployments.
@@
Finally, in the present study, the polarization sensitivity is essentially dictated by the adopted meta-atom geometry rather than by the cylindrical conformal arrangement itself. In particular, the asymmetric loading associated with the PIN diode breaks the unit-cell symmetry, resulting in a polarization-selective response. The cylindrical geometry mainly affects the local angle of incidence, while the polarization response remains governed by the meta-atom design. It is worth noting that polarization-insensitive metasurfaces have been widely reported (see, e.g., [24]). These approaches could be directly extended to cylindrical RISs to mitigate polarization sensitivity. A systematic investigation of polarization effects in conformal RISs, including cross-polar coupling induced by curvature, will be addressed in future work.